[RESOLVED] Broadband packet loss

We are currently investigating a packet loss issue affecting Broadband products only; ADSL, FTTC, FTTP. All other services are unaffected.
An upstream network supplier is investigating and we will update when we know more.

Posted in ADSL, Incidents
4 comments on “[RESOLVED] Broadband packet loss
  1. An upstream network provider has acknowledged the issue and identified Packet Loss across their core network. They have escalated the incident to their Network Engineers and a Technical call is ongoing to review how we stabilise service and alleviate any customer impact.

  2. The upstream suppliers NOC Engineers have escalated the incident to Senior Network Engineers to assist with investigations. At this juncture they have identified a potential capacity issue on their Core Network in London. A technical sessions is ongoing between Engineering Teams to determine a plan to move traffic across the network to alleviate any customer impact.

  3. The upstream network providers Technical Teams have completed configuration changes in attempt to mitigate customer impact, however, due to the drop in traffic volumes they are unable to confirm if this has had the desired effect. Testing confirms that Packet Loss has improved and the testing completed on the sample of Customer Premises Equipment (CPE) routers indicates no high levels of Packet Loss.

    A Technical Assistance Case (TAC) Case has been raised with Nokia to assist with their investigations to 1. replicate to issue and 2. provide guidance on best practice to distribute traffic across the core devices. Following the feedback from Nokia their Engineering Teams will review any recommendations and present any required changes via the Emergency Change Advisory Board (ECAB). They have gathered historical data to assist Nokia with their investigations.

    The initial increase in traffic today is thought to be as a result of Microsoft Patch Tuesday which may have caused an increase in traffic volumes, however, they are unable to confirm this until they see traffic levels increase on the 12th of May 2022.

    The incident has been placed into an extended monitoring status and in the event of a reoccurrence the Major Incident Team will reconvene.

    Next Update: 10:30am 12/05/2022

  4. Further configuration changes were made by the upstream network provider overnight. Our monitoring overnight and this morning indicates the issue is resolved.

Leave a Reply

Please log in using one of these methods to post your comment:

WordPress.com Logo

You are commenting using your WordPress.com account. Log Out /  Change )

Facebook photo

You are commenting using your Facebook account. Log Out /  Change )

Connecting to %s